Find the measure of an angle whose supplement and complement add to 194 degrees.
Answer by mananth(16949)

let angle be x
complement = 90-x
supplement = 180-x
..
180-x+90-x= 194
270-2x=194
2x=270-194
2x=76
x= 38 deg
...
m.ananth@hotmail.ca